WITHOUT graphing, tell whether each point is on the graph of the line given
12x+3y=6; (2,-4)
Please Show Steps, THANK YOU!!

Answer by nerdybill(7384)

When the problem gives you:
(2,-4)
which gives you the x and y value (x,y).
So, simply plug the values into the equation and check:
12x+3y=6
12(2)+3(-4)=6
24+(-12)=6
24-12=6
12=6 (FALSE)
since, the above is false, we conclude that
(2,-4) is NOT a solution for 12x+3y=6